By referring to the chemical equation given, calculate the volume of nitrogen gas in litres, required to react with 40.0 g of magnesium at a pressure of 2.25 atm and a temperature of 15 °C.

            3Mg (s) + N2 → Mg3N2 (s)

1
Expert's answer
2021-11-12T07:46:02-0500

3Mg + N2 → Mg3N2

M(Mg) = 24.3 g/mol

n(Mg) "= \frac{40.0}{24.3} = 1.646 \;mol"

According to the reaction equation

n(N2"= \frac{1}{3}n(Mg) = \frac{1}{3} \times 1.646 = 0.548 \;mol"

Ideal gas low

pV = nRT

p= 2.25 atm

T=15 + 273 = 288 \;K

R = 0.08206 L×atm/mol×K

"V = \frac{nRT}{p} \\

V (N_2) = \frac{0.548 \times 0.08206 \times 288}{2.25} = 5.756 \;L"

Answer: 5.756 L